OpenStudy (anonymous):

Simplifysquare root of 3 multiplied by the fifth root of 3. 3 to the power of 1 over 10 3 to the power of 3 over 5 3 to the power of 9 over 10 3 to the power of 7 over 10

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Hi, could you please help @phi

OpenStudy (anonymous):

I think it could be C

OpenStudy (phi):

can you write both numbers using exponents?

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

(square root 3) (*5 square root 3)

OpenStudy (phi):

You should learn that \[ \sqrt{x} = x^\frac{1}{2} \]

OpenStudy (anonymous):

I dont know it wont let me

OpenStudy (phi):

the fifth root of 3 means you can write 3 with an exponent of 1/5 follow ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So then what would I do with the other three, the one without the xponent?

OpenStudy (phi):

to answer the question write the numbers as 3^(1/2) * 3^(1/5)

OpenStudy (anonymous):

Thank you soo much!

OpenStudy (phi):

and use the rule \[ a^b \cdot a^c = a^{b+c} \]

OpenStudy (phi):

can you add 1/2 + 1/5 ? what do you get ?

OpenStudy (anonymous):

would it be 7/10?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

or .7

OpenStudy (phi):

yes. so the answer is 3^(7/10)

OpenStudy (anonymous):

Oh thank you so so much!

OpenStudy (phi):

yes 3^0.7 is also correct. But as you see they gave you the choice of 3^(7/10) which means the same thing.

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

so to answer, learn that sqrt(x) is the same as x^(1/2) and "fifth root" means x^(1/5) and when we multiply the same base , we can add the exponents

OpenStudy (anonymous):

Oh ok, I will keep that in mind